### v1

# Parametric All-Optical Modulation on Chip

Abstract
We demonstrate parametric all-optical modulation in a periodically-poled lithium niobate microring resonator on chip. It employs quantum Zeno blockade between two distinct waves, a signal and a pump, through their sum-frequency generation at a large per-photon efficiency of MHz. With nanosecond pump pulses at mW peak power, modulation extinction is observed, marking over times efficiency improvement across various previous implementations. With only mW pump peak power, modulation extinction is observed for a doubly-stronger signal at mW. This demonstrates, for the first time, that optical transistors with cascadability and fan-out are possible with just parametric nonlinear optics. These results, together with inherent advantages in such photonic integrated circuits, open the door to scalable technology for all-optical and quantum information processing.
Introduction— Quantum Zeno effect occurs when a coherently-evolving system is coupled strongly to external degrees of freedom, with the result that the evolution is suppressed or frozen [1, 2, 3, 4]. Since its first demonstration in trapped ions in 1990 [5], the Zeno effect has been widely studied as a counter-intuitive and intriguing quantum phenomenon [6, 7, 8, 9, 10], and explored for exotic applications in quantum communications [11], quantum computing [12], all-optical information processing [13, 14, 15], and so on. Among then, quantum Zeno blockade (QZB) was analyzed and demonstrated in nonlinear optical systems, including those of waveguides [16, 17], etalons with sandwiched crystals [18], whispering-gallary mode cavities [14], a thin-film microdisk [19], and aluminum nitride microrings [20].
Optical QZB occurs in an open quantum system consisted of certain optical modes coupled with external degrees of freedom through a nonlinear-optical channel. When the channel is “fast” (i.e., the nonlinear interaction is strong), the occupation of a mode can transiently alter the system condition and suppress its coherent dynamics, including “block” additional photons from coupling into a certain mode [21, 22, 13, 23]. In a cavity setting, QZB can be implemented through sum-frequency generation (SFG) or difference-frequency generation (DFG), where populating the cavity with pump photons frustrates the cavity resonance condition for signal photons at a different wavelength [22, 13, 18, 14]. As a result, the signal is deflected from the cavity, and the frequency generation does not occur in the asymptotic limit. During this process, the pump photons act effectively as a probe dynamically monitoring if signal photons have entered the cavity, by coupling the latter to another lossy mode [24]. When the monitoring is fast, the Zeno effect prevents the signal from the cavity, thus the blockade effect [1, 2, 3, 5].
Hence, QZB provides a counter-intuitive implementation of the interaction between optical signals: one to control another without them physically overlapping to interact [7, 8]. Furthermore, unlike the photon blockade mediated by single emitters, here no excitation of any kind is created, but the interaction occurs due to the mere potential for a parametric nonlinear optical process (which asymptotically does not occur) [16, 22]. These distinct interaction-free and excitation-free features together eliminate the occurrence of phase noise, photon dissipation, and quantum state decoherence (e.g., via spontaneous emission) [25]. It thus promises to overcome the fundamental challenges facing all-optical processing of quantum signals, including realizing deterministic interaction between single photons for scalable photonic quantum computing [24].
The same interaction-free and excitation-free implementation of interaction makes QZB is strong candidate for optical transistors [26], the backbone technology for of the future all-optical information processing. Asides from the above practical advantages in noise, loss, and decoherence, QZB naturally support superior cascadability, fan out, and absence of critical biasing, as critical for practical optical logic. This is because QZB effect transiently changes the cavity condition, so that the blockade effect is independent of the signal’s power level and detailed waveforms. It thus allows a weaker pump to control a strong signal, or the output of one stage to drive multiple subsequent stages. Notably, as QZB only utilizes nonlinear optics, there is no energy conversion from optics to electronics or atomic, so that the otherwise detrimental heat sinking or electrical interference issues are eliminated.
In view of those potentials, here we demonstrate, for the first time, a chip-integrated device for QZB and efficient all-optical modulation with cascadability and fan-out. Unlike previous QZB demonstrations, we use etched and encapsulated microring cavities on thin-film lithium niobate [27, 28]. By strong cavity confinement and quasi-phase matching, we observe high modulation extinction ratio (MER) requiring only milliwatt peak power, pico-joule pulse energy, corresponding to over orders of magnitude improvement over the state of the art (see Table 1 below). For cascadability and fan out, we also demonstrate that a weak pump can modulate a twice as strong signal. Together, our results highlight QZB in integrated lithium niobate nanophotonics as an enabling candidate for the future all-optical logic, photonic quantum computing, and beyond.
Reference | Pump Peak Power | MER |
[18] | W | |
[19] | mW | |
[20] | mW | |
This work | mW |
Experiments and simulations The periodically poled lithium niobate micro-ring (PPLNMR), whose input-output is through a pulley bus waveguide, was fabricated and characterized using procedures detailed in our previous works [29, 30, 31, 32]. We utilize the largest tensor of lithium niobate’s in 600 nm z-cut samples and fabricate PPLNMR chip with a photon-photon coupling coefficient of MHz (angular frequency) and an overall intrinsic Quality factor . It supports quasi-phase matched sum-frequency generation among three quasi-transverse magnetic (quasi-TM) modes, the signal at nm, the pump at nm, and their sum-frequency at nm. Their intrinsic decay rates due to material absorption and scattering losses, external decay rates due to coupling with the pulley waveguide, and the total rates, are measured and listed in Table 2. As shown, for all three waves, the external decay is faster than the intrinsic decay by at least four times, which puts this chip in the significant over-coupling regime, as needed for most optical information processing.
Intrinsic Decay | External Decay | Total Decay | |
Pump | GHz | GHz | GHz |
Signal | GHz | GHz | GHz |
Sum-frequency | GHz | GHz | GHz |
Note that those measurements in Table 2 were performed at very low light intensity level, when the cavity is “cold” (room temperature). During experiment, the cavity may be heated up under high light intensity, and those parameters may change modestly. Meanwhile, the phase matching condition will change, which will require small change in chip temperature and the Pump and Signal wavelength to restore.
The cavity dynamics can be modeled by the following equations [33]:
In the above equations, the subscripts denote the pump, signal, and sum-frequency waves, respectively. is the external wave at the input. is the amplitude of the cavity field normalized such that gives the photon number in the cavity. is the cavity output. The total cavity decay rate is the sum of the intrinsic decay due to cavity losses and the external decay through coupling with the bus waveguide. is the laser-cavity detuning, which is zero under ideal phase matching. With cold cavity , the simulation results are calculated numerically.
The experiment setup is depicted in the Fig.1. It consists of two narrow-band tunable lasers (Santec TSL-550/570), each followed by electronics and optics for generating synchronized optical pulse. They include an arbitrary waveform generator (AWG) generating synchronous electrical pulses to drive electric-optical modulators (EOMs) at MHz for the pump light and MHz for the signal light. The resulting optical pulse trains each undergo amplification by an Erbium-doped fiber amplifier (EDFA), followed by wavelength division multiplexers (WDMs) to filter out excessive ASE noise. Then, attenuators (ATTs) are used to vary each optical power level, before combining via a beamsplitter and directed jointly into the PPLNMR chip via a lensed fiber. The chip sits on a thermoelectric cooler (TEC) that controls and stabilizes its temperature with precision. After the chip, another WDM is used to separate the output pump and signal, before each is detected by a photodetector and an oscilloscope (HP Agilent 54750a).
Continuous-wave Signal and Pulsed Pump— We first demonstrate all-optical modulation where a continuous-wave (CW) signal is modulated by a pulsed pump. To minimize thermal and photo-refractive (PR) effects [34][35][36], the signal is prepared in a rectangular waveform with -ns quasi-Continuous Wave (quasi-CW) pulses at MHz repetition rate and mW peak power on chip (note that the power number to be quoted for the rest of the paper always refers to the peak power on chip). The pump pulses, on the other hand, are in Gaussian form with -ns full width half maximum (FWHM), at MHz repetition rate, and are set in the middle of the signal pulses.
The results are shown in Fig. 2(a-d). When there is no pump, the on-resonance signal transmission loss through the chip is as compared to the off resonance case, due to the signal over-coupling. Then, as shown in Fig. 2(b), when pump pulses with low optical power ( mW) are applied, a dip is created in their overlapping regime, with the transmission dropping to nearly zero (e.g., 100% loss). This is because the pump drives SFG in the cavity, inducing additional losses and causing the cavity to become critically coupled for the signal.
In Fig. 2(c), as the pump power gradually increases to mW, the signal transmission bounces back. At the peak of the pump pulses, the transmission loss is reduced to , which is about half of the on-resonance case. This is the result of quantum Zeno blockade: when the SFG-induced cavity loss is larger than the coupling decay, the signal is effectively under coupled from the cavity, so that it mostly does not enter the cavity [16, 18]. The MER, calculated as the percentage reduction of transmission due to the pump, is . As shown in Fig. 2(d), further increasing the pump power to mW improves the MER to with even lower transmission loss . The trenches flanking the peak indicates effective critical coupling for the signal in those regimes, where the pump power induced cavity loses for the signal through SFG but not strong enough to cause QZB. In both Figs. 2(c) and 2(d), the SFG at the pump peak is suppressed compared to the weaker pump case in Fig. 2(b). Again, this is because QZB blocks the signal from entering the cavity.
Also plotted in Figs. 2(a-d) are the simulation results by solving Eqs.(1-6) with only the separately measured cold-cavity parameters in Table 2, without using any fitting parameter. As shown, the theoretical results follow closely the experimental results, indicating a good modeling of underlying dynamics. The disagreement in Figs. 2(c) and (d) may due to the an over-estimate of the SFG efficiency and slight shift of the signal resonance peak in a “hot” cavity with high pump power. Note that in Figs. 2(c-d), the transmitted signals have some oscillations around the peaks. This could be due to the actual coupling condition for the SF wave is less over-coupled, so that a strong SF field is trapped in the cavity, causing transient parametric oscillation. More discussions can be found in the Supplemental Materials.
Pulsed Signal and Pulsed Pump— Next we demonstrate the modulation of signal pulses by pump pulses. The results for Gaussian signal pulses at -ns FWHM and with the same peak power are shown in Fig.2(e-h). With a much lower average power for the signal, the chip temperature and thus the cavity-coupling condition are changed, giving a on-resonance transmission loss (which indicates deeper over-coupling). As seen, with the same pump applied, the modulation results follow the trend with the above quasi-CW case. That’s when the signal overlapped with 0.25-mW pump pulses, the signal undergoes further suppression to critical coupling. Increasing the pump power to mW results in the signal transmission loss lowered to and MER reaching .
The simulation results plotted in Fig.2(e-h), using the measured pulse profiles and amplitudes and cold-cavity parameters in Table 2, give a signal transmission dynamic mostly aligned with the experiment results.
Strong Signal and Weak Pump— When the power of pump pulses reaches the threshold to frustrate the cavity, they can modulate and switch signals of much higher power. This unique feature, originating from QZB, is essential for scalability, fan out, and absence of critical bias in practical optical transistors. In Fig. 3, we demonstrate weaker pump pulses modulating stronger signal pulses. The signal and pump pulse shapes are the same with Fig. 2, but with the pump at mW and the signal at mW. As shown, the signal on resonance transmission loss is . By adjusting the phase on AWG, the signal arrival time is tuned to achieve the best MER of at a ps delay from the pump. Its transmission loss is reduced to .
Figure 3 also plots the simulation results by using measured pulse profiles, amplitudes, and the cold-cavity parameters, without any fitting parameters. As seen, they are in good agreement with the experimental results. The on-resonance transmission loss is and the MER is , with transmission loss under the pump modulation. Compared with the experiment results, the simulated MER is slightly lower. This is attributed to that the actual signal coupling is weaker than the measured cold-cavity case. Also, the simulated signal transmission shows a ringing bell oscillation. This is because at the pump pulse tail, SFG is not strong enough to prohibit the signal from entering the cavity via QZB, but rather causing critical coupling for the signal. Further into the tail, the SFG is diminished, because by then the pump has exited the cavity. Thus the signal “sees” an overcoupling cavity, so that its transmission loss is less. Yet this ringing is not observed in experiment, which indicates other effects, such as those of photorefractive (PR), thermo-optic, and second-harmonic generation (SHG), may need to be added to the current model for a more accurate description of the system dynamics. This will remain a task for the future studies.
Conclusion and Discussion— We have demonstrated parametric all-optical modulation with high efficiency and presented clear evidences towards its uses for practical optical transistor technology. It employs quantum Zeno blockade in a periodic poled microring in thin-film lithium niobate, offering rich and counter-intuitive nonlinear optical phenomena. Our simulation results agree well with the experimental results, with subtle differences coming from side effects not captured by our model, including those of PR, thermo-optic, and SHG. They call for suppression of those effects for improved performance, e.g., by addressing PR by MgO doping and addressing SHG by narrowing phase matching. To reduce the thermo-optic effect, one can continuously monitor the signal at an adjacent period to make sure it stays on resonance and fine adjustment of the phase matching condition by tuning the tunable lasers and the TEC.
From our results, the modulation performance is not sensitive to the pulse length, as long as it exceeds the cavity life-time. In fact, in the “Pulsed Signal and Pulsed Pump” case, a higher MER of is seen for the same pump power than in the case with a quasi-CW signal. Also, the MER can be increased by operating in a less over-coupling regime. As of now, the power consumption for modulation is 12 pJ. By using shorter pulses closer to the cavity lifetime of ps, the modulation energy can be reducted to pJ, although we currently do not have this instrument to create such pulses.
We would like to note that the idea of light controlling light has gained interest in the pursuit of all-optical information processing and beyond, as an indispensable element and a powerful tool. Recent demonstrations have employed two-photon absorption [37] and second-harmonic generation and its reversion [38]. In comparison, QZB uniquely allows all-optical logic in an interaction-free and excitation-free manner, promising chip-integrated optical transistor technology that is robust and scalable. For example, in the Supplemental Materials, we show how an all-optical switch is realized simply by adding a drop port. For realistic device parameters and 1 ns pulses, a 6-mW signal can be switched by a 2-mW pump with over 90% efficiency. Meanwhile, due to interaction-free, of pump survives after the switch, so that it can subsequently switch another signal pulse.
Another interesting feature with QZB is that, the roles of control and signal are determined by which arrives at the cavity earlier. This gives rise to a potentially powerful operation where the conditional logic is reversed by controlling the relative time delay between the input pulses. For detailed result figures please see the Supplemental Materials .
Lastly, the interaction-free and excitation-free features make QZB particularly suitable for quantum applications. Indeed, the same device shall readily work for quantum signals at a single-photon level, except for potential Raman scattering which can be suppressed by further detuning the pump and signal. Furthermore, by improving the intrinsic Q-factor to , as demonstrated in a laser-written and chemical and mechanical polished microring [39], the required pump power can be reduced to single-photon level as well. This opens door to single-photon anharmonicity and quantum C-NOT gates between single photons.
Acknowledgements.
The research was supported in part by the Office of Naval Research (Award No. N00014-21-1-2898). Device fabrication was performed at Advanced Science Research Center, City University of New York.References
- Misra and Sudarshan [1977] B. Misra and E. G. Sudarshan, The zeno’s paradox in quantum theory, Journal of Mathematical Physics 18, 756 (1977).
- Peres [1980] A. Peres, Zeno paradox in quantum theory, American Journal of Physics 48, 931 (1980).
- Joos [1984] E. Joos, Continuous measurement: Watchdog effect versus golden rule, Physical Review D 29, 1626 (1984).
- Ai et al. [2010] Q. Ai, Y. Li, H. Zheng, C. Sun, et al., Quantum anti-zeno effect without rotating wave approximation, Physical Review A 81, 042116 (2010).
- Itano et al. [1990] W. M. Itano, D. J. Heinzen, J. J. Bollinger, and D. J. Wineland, Quantum zeno effect, Physical Review A 41, 2295 (1990).
- Elitzur and Vaidman [1993] A. C. Elitzur and L. Vaidman, Quantum mechanical interaction-free measurements, Foundations of physics 23, 987 (1993).
- Kwiat et al. [1995] P. Kwiat, H. Weinfurter, T. Herzog, A. Zeilinger, and M. A. Kasevich, Interaction-free measurement, Physical Review Letters 74, 4763 (1995).
- Kwiat et al. [1996] P. Kwiat, H. Weinfurter, and A. Zeilinger, Quantum seeing in the dark, Scientific American 275, 72 (1996).
- Tsegaye et al. [1998] T. Tsegaye, E. Goobar, A. Karlsson, G. Björk, M. Loh, and K. Lim, Efficient interaction-free measurements in a high-finesse interferometer, Physical Review A 57, 3987 (1998).
- Kwiat et al. [1999] P. G. Kwiat, A. White, J. Mitchell, O. Nairz, G. Weihs, H. Weinfurter, and A. Zeilinger, High-efficiency quantum interrogation measurements via the quantum zeno effect, Physical Review Letters 83, 4725 (1999).
- Cao et al. [2017] Y. Cao, Y.-H. Li, Z. Cao, J. Yin, Y.-A. Chen, H.-L. Yin, T.-Y. Chen, X. Ma, C.-Z. Peng, and J.-W. Pan, Direct counterfactual communication via quantum zeno effect, Proceedings of the National Academy of Sciences 114, 4920 (2017).
- Hosten et al. [2006] O. Hosten, M. T. Rakher, J. T. Barreiro, N. A. Peters, and P. G. Kwiat, Counterfactual quantum computation through quantum interrogation, Nature 439, 949 (2006).
- Huang and Kumar [2011] Y.-P. Huang and P. Kumar, Interaction-free quantum optical fredkin gates in microdisks, IEEE Journal of Selected Topics in Quantum Electronics 18, 600 (2011).
- Strekalov et al. [2014] D. V. Strekalov, A. S. Kowligy, Y.-P. Huang, and P. Kumar, Progress towards interaction-free all-optical devices, Physical Review A 89, 063820 (2014).
- Chang et al. [2014] D. E. Chang, V. Vuletić, and M. D. Lukin, Quantum nonlinear optics—photon by photon, Nature Photonics 8, 685 (2014).
- Huang et al. [2010] Y.-P. Huang, J. B. Altepeter, and P. Kumar, Interaction-free all-optical switching via the quantum zeno effect, Physical Review A 82, 063826 (2010).
- Liu et al. [2023] Q. Liu, W. Liu, K. Ziegler, and F. Chen, Engineering of zeno dynamics in integrated photonics, Physical Review Letters 130, 103801 (2023).
- McCusker et al. [2013] K. T. McCusker, Y.-P. Huang, A. S. Kowligy, and P. Kumar, Experimental demonstration of interaction-free all-optical switching via the quantum zeno effect, Physical review letters 110, 240403 (2013).
- Chen et al. [2017] J.-Y. Chen, Y. M. Sua, Z.-T. Zhao, M. Li, and Y.-P. Huang, Observation of quantum zeno blockade on chip, Scientific reports 7, 14831 (2017).
- Guo et al. [2018] X. Guo, C.-L. Zou, L. Jiang, and H. X. Tang, All-optical control of linear and nonlinear energy transfer via the zeno effect, Physical Review Letters 120, 203902 (2018).
- Jacobs and Franson [2009] B. C. Jacobs and J. Franson, All-optical switching using the quantum zeno effect and two-photon absorption, Physical Review A 79, 063830 (2009).
- Huang and Kumar [2010] Y.-P. Huang and P. Kumar, Interaction-free all-optical switching in (2) microdisks for quantum applications, Optics letters 35, 2376 (2010).
- Huang and Kumar [2012] Y.-P. Huang and P. Kumar, Antibunched emission of photon pairs via quantum zeno blockade, Physical review letters 108, 030502 (2012).
- Sun et al. [2013] Y.-Z. Sun, Y.-P. Huang, and P. Kumar, Photonic nonlinearities via quantum zeno blockade, Physical review letters 110, 223901 (2013).
- Shapiro [2006] J. H. Shapiro, Single-photon kerr nonlinearities do not help quantum computation, Physical Review A 73, 062305 (2006).
- Miller [2010] D. A. Miller, Are optical transistors the logical next step?, Nature Photonics 4, 3 (2010).
- Zhu et al. [2021] D. Zhu, L. Shao, M. Yu, R. Cheng, B. Desiatov, C. Xin, Y. Hu, J. Holzgrafe, S. Ghosh, A. Shams-Ansari, et al., Integrated photonics on thin-film lithium niobate, Advances in Optics and Photonics 13, 242 (2021).
- Lin et al. [2020] J. Lin, F. Bo, Y. Cheng, and J. Xu, Advances in on-chip photonic devices based on lithium niobate on insulator, Photonics Research 8, 1910 (2020).
- Chen et al. [2019] J.-Y. Chen, Z.-H. Ma, Y. M. Sua, Z. Li, C. Tang, and Y.-P. Huang, Ultra-efficient frequency conversion in quasi-phase-matched lithium niobate microrings, Optica 6, 1244 (2019).
- Ma et al. [2020] Z. Ma, J.-Y. Chen, Z. Li, C. Tang, Y. M. Sua, H. Fan, and Y.-P. Huang, Ultrabright quantum photon sources on chip, Physical Review Letters 125, 263602 (2020).
- Chen et al. [2021] J.-Y. Chen, Z. Li, Z. Ma, C. Tang, H. Fan, Y. M. Sua, and Y.-P. Huang, Photon conversion and interaction in a quasi-phase-matched microresonator, Physical Review Applied 16, 064004 (2021).
- Ma et al. [2023] Z. Ma, J.-Y. Chen, M. Garikapati, Z. Li, C. Tang, Y. M. Sua, and Y.-P. Huang, Highly efficient and pure few-photon source on chip, Physical Review Applied 20, 044033 (2023).
- Chin and Ho [1998] M. Chin and S. Ho, Design and modeling of waveguide-coupled single-mode microring resonators, Journal of lightwave technology 16, 1433 (1998).
- Chen et al. [2020] J.-Y. Chen, C. Tang, Z.-H. Ma, Z. Li, Y. M. Sua, and Y.-P. Huang, Efficient and highly tunable second-harmonic generation in z-cut periodically poled lithium niobate nanowaveguides, Optics Letters 45, 3789 (2020).
- Surya et al. [2021] J. B. Surya, J. Lu, Y. Xu, and H. X. Tang, Stable tuning of photorefractive microcavities using an auxiliary laser, Optics Letters 46, 328 (2021).
- Shams-Ansari et al. [2022] A. Shams-Ansari, G. Huang, L. He, Z. Li, J. Holzgrafe, M. Jankowski, M. Churaev, P. Kharel, R. Cheng, D. Zhu, et al., Reduced material loss in thin-film lithium niobate waveguides, Apl Photonics 7 (2022).
- Grinblat et al. [2019] G. Grinblat, M. P. Nielsen, P. Dichtl, Y. Li, R. F. Oulton, and S. A. Maier, Ultrafast sub–30-fs all-optical switching based on gallium phosphide, Science advances 5, eaaw3262 (2019).
- Guo et al. [2022] Q. Guo, R. Sekine, L. Ledezma, R. Nehra, D. J. Dean, A. Roy, R. M. Gray, S. Jahani, and A. Marandi, Femtojoule femtosecond all-optical switching in lithium niobate nanophotonics, Nature Photonics 16, 625 (2022).
- Gao et al. [2022] R. Gao, N. Yao, J. Guan, L. Deng, J. Lin, M. Wang, L. Qiao, W. Fang, and Y. Cheng, Lithium niobate microring with ultra-high q factor above 10 8, Chinese Optics Letters 20, 011902 (2022).